OpenStudy (anonymous):

MEDAL simplify step by step please: 2-ln(x)^2

OpenStudy (ipwnbunnies):

\[2-\ln(x^{2})?\]

OpenStudy (ipwnbunnies):

This is the power property of logs: \[\log(x^{a}) = a*\log(x)\]

OpenStudy (ipwnbunnies):

log can also be ln.